# Transparent and Coherent Procedural Mistake Detection

### v2

@@ Abstract
Procedural mistake detection (PMD) is a challenging problem of classifying whether a human user (observed through egocentric video) has successfully executed a task (specified by a procedural text). Despite significant recent efforts, machine performance in the wild remains nonviable, and the reasoning processes underlying this performance are opaque. As such, we extend PMD to require generating visual self-dialog rationales to inform decisions. Given the impressive, mature image understanding capabilities observed in recent vision-and-language models (VLMs), we curate a suitable benchmark dataset for PMD based on individual frames. As our reformulation enables unprecedented transparency, we leverage a natural language inference (NLI) model to formulate two automated metrics for the coherence of generated rationales. We establish baselines for this reframed task, showing that while VLMs struggle off-the-shelf, their accuracy, coherence, and efficiency can be improved by incorporating these metrics into common inference and fine-tuning methods- though not without tradeoff. Lastly, our multi-faceted metrics visualize common outcomes, highlighting areas for further improvement.
@@ 1 Introduction
The problem of interactive task guidance has recently attracted attention in the AI research community Bao et al. (2023); Wang et al. (2023a); Peddi et al. (2024); Bohus et al. (2024), stemming from significant efforts to build and learn from large-scale procedural video datasets Zhou et al. (2018); Damen et al. (2018); Miech et al. (2019); Grauman et al. (2022). A successful task guidance agent can observe a human user through video and interact with them through language to guide them to complete a task. A key component of such an agent is procedural mistake detection (PMD): the ability to detect when the user’s actions deviate from a procedural text, e.g., a recipe or instruction manual. To achieve this, a system must anticipate the success conditions for actions in this procedural text, then verify them by extracting relevant physical state information from the visual scene.
@@
Prior work in PMD has explored a mix of specialized (primarily vision-based) classifiers Sener et al. (2022); Wang et al. (2023a); Peddi et al. (2024) as well as foundational language models (LMs) and vision-and-language models (VLMs) Du et al. (2023); Bao et al. (2023); Peddi et al. (2024); Flaborea et al. (2024), but this problem has proven difficult, and current approaches fail to achieve accuracy above chance in zero-shot PMD in the wild. Qualitatively, Bao et al. (2023) find that while the web-scale multimodal pre-training of foundational VLMs enables coverage of a wide variety of procedures, they often produce noisy, vague, or otherwise insufficient information from visual scenes to facilitate PMD. This capability to extract and reason over key task-relevant visual information is crucial to PMD, but prior work has largely overlooked it, instead targeting binary or categorical classification tasks in their system design and quantitative evaluations. Consequently, the reasons for VLMs’ decisions are opaque, hindering practical use [^1] Since VLMs are unreliable mistake detectors, transparency is crucial for the user to understand the system’s reasoning process and act on or disregard it accordingly. and continued improvement.
@@
To promote the development of PMD systems with transparent and justified decisions, we propose a reformulated problem of coherent PMD: given a procedural text and egocentric video frame, VLMs must not only classify whether a mistake has occurred, but also support this decision with a rationale consisting of evidence from the visual scene. As shown in Figure 1, this rationale takes the form of an iterative self-dialog of generated questions and yes-no answers which directly condition classification. [^2] This choice of rationale resembles visual question answering (VQA; Antol et al., 2015) and visual dialog Das et al. (2017), long-studied multimodal tasks that VLMs excel at Dai et al. (2023a); Liu et al. (2023); Dubey et al. (2024). Since recent VLMs struggle to extract detailed information from videos, but have exhibited more mature image understanding capabilities, we curate an approachable large-scale dataset for PMD based on individual video frames annotated in Ego4D Grauman et al. (2022). We define two metrics for the coherence of generated rationales based on a natural language inference (NLI) model. To lay a foundation for research in coherent PMD, we establish baselines by exploring three natural interventions to VLMs: (1) we use our metrics to re-rank candidate questions generated by VLMs, (2) we harness VLMs’ in-context learning capability to generate additional candidate questions based on human-written examples, and (3) we use our metrics to fine-tune VLMs to generate more coherent questions. Our results show that while VLMs struggle off-the-shelf, these interventions can improve VLMs’ accuracy, coherence, and rationale generation efficiency, albeit creating tradeoffs between these properties. We lastly show how our multi-faceted metrics visualize common outcomes in coherent PMD, enabling fine-grained evaluation and identification of areas for future improvement.
@@ 2.2 Applying VLMs to Coherent PMD
To prevent over-generating evidence, which could introduce noise and degrade PMD accuracy, we implement an early stopping mechanism to determine whether to stop generating questions based on the success likelihood after each iteration. The self-dialog stops early (i.e., before $n$ questions are generated) if the success likelihood stabilizes (i.e., changes by less than $\delta$ for two consecutive iterations) or becomes highly confident (i.e., subceeds $\epsilon$ or exceeds $1-\epsilon$). $n$ is fixed at 10, while $\delta$ and $\epsilon$ are tuned based on the validation data for each presented approach.
@@ 2.3 Constructing a Dataset for PMD
We follow Du et al. (2023) in recasting Ego4D Grauman et al. (2022), a large-scale egocentric video dataset for everyday activities with dense annotations for various aspects of the videos, into an offline mistake detection format, but expand the diversity of mistake types studied there. Ego4D’s hand and object interactions data subset includes videos of physical actions being performed with various objects. Each video is annotated with narrations describing fine-grained procedures being performed, timestamps for when it begins and ends, and category labels for the verb and noun characterizing the procedure. This makes an ideal testbed for evaluating VLMs’ understanding of real-world actions in video frames, but the data is not formulated for mistake detection. We thus apply several preprocessing steps to the data to create a new Ego4D for Procedural Mistake Detection (Ego4D-PMD) benchmark that includes successful cases and a breadth of mistake types for each annotated procedure, visualized in Figure 2.
Additional details about these preprocessing steps and summary statistics for the Ego4D-PMD dataset are presented in Appendix A. As listed there, to conserve compute, we randomly sample a subset of 10,000, 500, and 2,000 examples respectively from the training, validation, and testing partitions (evenly split between success and mistake cases) for the forthcoming experiments.
@@ 3.1 Using NLI Models to Judge Success
We implement $f_{e}$ with BART Lewis et al. (2020) fine-tuned on the large-scale MultiNLI dataset Williams et al. (2017), [^4] See https://huggingface.co/facebook/bart-large-mnli. applying softmax over its logits for entailment and contradiction to get an entailment probability. We follow Srinivasan et al. (2024) and prompt a foundational LM to implement $t_{p}$. [^5] To conserve GPU memory, we later use the evaluated VLM’s LM backbone for rephrasing. Details in Appendix B. For the procedural text $T$, we choose a success prompt template $t_{h}$ “The procedure $\langle P\rangle$ has been successfully executed.” [^6] A template was used as we found complex procedural texts were unlikely to be rephrased accurately, degrading correlation with human judgments (see Appendix C).
@@ 4.2 Coherent Question Generation
While the above approaches may boost the coherence of PMD rationales, they have limitations. First, since they are training-free, they do not improve the internal coherence of VLMs, rather they only filter and augment their outputs. Further, these interventions take significantly more time and compute due to the need to evaluate the coherence of candidate questions (which requires prompting an NLI model) and generate questions twice (once based on the self-dialog history and once with in-context learning). In practical applications like task guidance, it may be advantageous to use specialized VLMs to improve reliability and speed. As such, we also explore whether VLMs can be fine-tuned to generate more coherent questions using our automated coherence metrics.
@@ 5 Experimental Results
We specifically evaluate InstructBLIP Dai et al. (2023a), LLaVA 1.5-7B Liu et al. (2023), and Llama 3.2-Vision-11B Dubey et al. (2024), small open-source VLMs feasible for online use (important for real-world applications like task guidance). They apply different architectures and training strategies to integrate vision into their LMs. LLaVA and InstructBLIP were not trained on Ego4D, while Llama 3’s training data was not disclosed. For a reference point with state-of-the-art proprietary VLMs, we additionally include results with off-the-shelf GPT-4o OpenAI et al. (2024b). Additional details and prompt templates can be found in Appendix D and E.4.131313In Appendix E.5, we include an additional evaluation of VLMs without generating rationales (as done in prior work).
@@
Table 1: Ego4D-PMD test set results for GPT-4o OpenAI et al. (2024b) and likelihood-based (L) and coherence-based (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
@@ 5.1 Coherent Question Selection Results
Experimental results for coherent question selection interventions are presented in Table 2.141414Hyperparameters and validation results in Appendix E.6. We find that coherence-based re-ranking and in-context learning [^15] We visualize the distribution of question sources (i.e., in-context learning vs. full self-dialog history) in Appendix E.7. both sharply improve the relevance and informativeness in all models, reaching a respective 75.5% and 0.464 bits in LLaVA. This suggests that questions that VLMs find most likely are not naturally the most coherent. Interestingly, accuracy also jumps sharply for InstructBLIP and LLaVA to a maximum of 67.8%, demonstrating that coherent rationales are valuable to accurate PMD. The information gain under our question selection interventions is consistently higher, reaching a maximum of 0.663 bits in LLaVA. This suggests that VLMs can make more confident conclusions given more coherent rationales. [^16] To better understand performance gains here, in Appendix E.8, we contextualize these results with a naïve semantic diversity-based ranking baseline. Meanwhile, our best configurations outperform GPT-4o in accuracy, relevance, and informativeness at less than 10% of its size Abacha et al. (2025). While GPT-4o takes fewer iterations and exhibits higher information gain (thus making faster and more confident decisions), its rigidity as a closed proprietary VLM prevents further improvements.
@@ 5.2 Coherent Question Generation Results
However, we find that accuracy and informativeness drop slightly from the best results in Table 2 to a respective maximum of 64.2% and 0.340 bits. This suggests that asking more relevant questions is not enough to improve performance globally. We suspect this is due to an implicit difficulty trade-off of asking highly relevant questions, which strongly indicate success if answered one way, otherwise a mistake. For elaborate procedures, highly relevant questions may cover multiple aspects or states of a scene. As such, answering these more complex questions may be difficult for VLMs. For example, while the VLM-generated question “Is the soil placed around the seedling with the trowel in the person’s hand?” covers the success conditions of the procedure “Put some soil around the tomato seedling with the gardening trowel in your hand,” answering it requires understanding spatial relations between several objects (i.e., soil, seedling, trowel, and hand). Future work may explore methods to encourage VLMs to generate simpler questions. [^17] As an initial inquiry here, we experiment with applying an exponential length penalty to VQG in Appendix E.9.
@@ 5.3 Visualizing PMD Performance
An additional benefit of our automated coherence metrics is the ability to audit the reasoning behaviors of VLMs. In Figure 4, we visualize the distribution of decision error, example-level relevance, and example-level informativeness of four representative approaches. Specifically, we compare vanilla LLaVA (i.e., with likelihood-based question ranking) to variants successively equipped with coherence-based question ranking, in-context learning, and coherence-based fine-tuning with DPO. For each example, decision error is calculated by how far the VLM’s success likelihood is from being 100% confident in the ground truth success decision. Point colors indicate combinations of decision error, relevance, and informativeness, highlighting common local outcomes. Several examples are described in Appendix E.10.

### v3

@@ Abstract
Procedural mistake detection (PMD) is a challenging problem of classifying whether a human user (observed through egocentric video) has successfully executed a task (specified by a procedural text). Despite significant recent efforts, machine performance in the wild remains nonviable, and the reasoning processes underlying this performance are opaque. As such, we extend PMD to require generating visual self-dialog rationales to inform decisions. Given the impressive, mature image understanding capabilities observed in recent vision-and-language models (VLMs), we curate a suitable benchmark dataset for PMD based on individual frames. As our reformulation enables unprecedented transparency, we leverage a natural language inference (NLI) model to formulate two automated metrics for the coherence of generated rationales. We establish baselines for this reframed task, showing that VLMs struggle off-the-shelf, but with some trade-offs, their accuracy, coherence, and efficiency can be improved by incorporating these metrics into common inference and fine-tuning methods. Lastly, our multi-faceted metrics visualize common outcomes, highlighting areas for further improvement.
@@ 1 Introduction
The problem of interactive task guidance has recently attracted attention in AI research Bao et al. (2023); Wang et al. (2023a); Peddi et al. (2024); Bohus et al. (2024), stemming from significant efforts to build and learn from large-scale procedural video datasets Zhou et al. (2018); Damen et al. (2018); Miech et al. (2019); Grauman et al. (2022). A successful task guidance agent can observe a human user through video and guide them to complete a task through language interaction. A key component of such an agent is procedural mistake detection (PMD): the ability to detect when the user’s actions deviate from a procedural text, e.g., a recipe or instruction manual. To achieve this, a system must apply physical and procedural commonsense knowledge to anticipate success conditions for the text, then extract relevant state information from the visual scene to verify them.
@@
Prior work in PMD has explored a mix of specialized (primarily vision-based) classifiers Sener et al. (2022); Wang et al. (2023a); Peddi et al. (2024) as well as foundational language models (LMs) and vision-and-language models (VLMs) Du et al. (2023); Bao et al. (2023); Peddi et al. (2024); Flaborea et al. (2024), but this problem has proven difficult, and current approaches fail to achieve accuracy above chance in zero-shot PMD in the wild. Qualitatively, Bao et al. (2023) found that while the web-scale multimodal pre-training of foundational VLMs enables coverage of a wide variety of procedures, they often produce noisy, vague, or otherwise insufficient information from visual scenes to facilitate PMD. This capability to extract and reason over key task-relevant visual information is crucial to PMD, but prior work has largely overlooked it, instead targeting binary or categorical classification tasks in their system design and quantitative evaluations. Consequently, the reasons for VLMs’ decisions are opaque, hindering practical use [^1] Since VLMs are unreliable mistake detectors, transparency is crucial for the user to understand the system’s reasoning process and act on or disregard it accordingly. and continued improvement.
@@
To promote the development of PMD systems with transparent and justified decisions, we propose a reformulated problem of coherent PMD: given a procedural text and egocentric video frame, VLMs must not only classify whether a mistake has occurred, but also support this decision with a rationale consisting of evidence from the visual scene. As shown in Figure 1, this rationale takes the form of an iterative self-dialog of generated questions and yes-no answers which directly condition classification. [^2] This choice of rationale resembles visual question answering (VQA; Antol et al., 2015) and visual dialog Das et al. (2017), long-studied multimodal tasks that VLMs excel at Dai et al. (2023a); Liu et al. (2023); Dubey et al. (2024). Since recent VLMs struggle to extract detailed, temporally coherent information from videos, but have exhibited more mature image understanding capabilities, we curate an approachable large-scale dataset for PMD based on individual video frames annotated in Ego4D Grauman et al. (2022). We define two metrics for the coherence of generated rationales based on a natural language inference (NLI) model. To lay a foundation for research in coherent PMD, we establish baselines by exploring three natural interventions to VLMs: (1) we use our metrics to re-rank candidate questions generated by VLMs, (2) we harness VLMs’ in-context learning capability to generate additional candidate questions based on human-written examples, and (3) we use our metrics to fine-tune VLMs to generate more coherent questions. Our results show that while VLMs struggle off-the-shelf, these interventions can improve VLMs’ accuracy, coherence, and rationale generation efficiency, albeit creating tradeoffs between these aspects. We lastly show how our multi-faceted metrics visualize common outcomes in coherent PMD (e.g., unjustified decisions, object hallucination, and more), enabling fine-grained evaluation and identification of areas for future improvement.
@@ 2.2 Applying VLMs to Coherent PMD
To prevent over-generating evidence, which could introduce noise and degrade PMD accuracy, we implement an early stopping mechanism to determine whether to stop generating questions based on the success likelihood after each iteration. The self-dialog stops early (i.e., before a maximum number of questions $n^{*}$ are generated) if the success likelihood stabilizes (i.e., changes by less than $\delta$ for two consecutive iterations) or becomes highly confident (i.e., subceeds $\epsilon$ or exceeds $1-\epsilon$). $n^{*}$ is fixed at 10, while $\delta$ and $\epsilon$ are tuned based on the validation data for each presented approach.
@@ 2.3 Constructing a Dataset for PMD
We follow Du et al. (2023) in recasting Ego4D Grauman et al. (2022), a large-scale egocentric video dataset for everyday activities with dense annotations for various aspects of the videos, into an offline mistake detection format, but expand the diversity of mistake types studied there. Ego4D’s hand and object interactions data subset includes videos of physical actions being performed with various objects. Each video is annotated with narrations describing fine-grained procedures being performed, each with timestamps for when it begins and ends, and category labels for the verb and noun characterizing the procedure. This makes an ideal testbed for evaluating VLMs’ understanding of real-world actions in video frames, but the data is not formulated for PMD. We thus apply several preprocessing steps to the data to create a new Ego4D for Procedural Mistake Detection (Ego4D-PMD) benchmark that includes successful cases and a breadth of mistake types for each annotated procedure, visualized in Figure 2.
Specifically, we form a successful case from each video clip in Ego4D by pairing its annotated postcondition frame with its annotated natural language narration of the procedure converted into imperative form. For each successful example, we generate several types of mistake examples paired with the source video’s procedural text. To simulate the procedure being incomplete, we follow Du et al. (2023) and sample another frame at the procedure’s precondition time. To simulate an incorrect action being applied and/or the action being applied to an incorrect object or ingredient, we sample alternative video clips with mismatched verbs and nouns in their narration texts. Additional details about these preprocessing steps and summary statistics for the Ego4D-PMD dataset are presented in Appendix A. As listed there, to conserve compute, we randomly sample a subset of 10,000, 500, and 2,000 examples respectively from the training, validation, and testing partitions (evenly split between success and mistake cases) for the forthcoming experiments.
@@ 3.1 Using NLI Models to Judge Success
We implement $f_{e}$ with BART Lewis et al. (2020) fine-tuned on the large-scale MultiNLI dataset Williams et al. (2017), [^4] See https://huggingface.co/facebook/bart-large-mnli. applying softmax over its logits for entailment and contradiction to get an entailment probability. We follow Srinivasan et al. (2024) and prompt a foundational LM to implement $t_{p}$. [^5] To conserve GPU memory, we later use the evaluated VLM’s LM backbone for rephrasing. Details in Appendix B. For the procedural text $T$, we choose a success prompt template $t_{h}$ “The procedure $\langle P\rangle$ has been successfully executed.” [^6] A template was used as we found complex procedural texts were unlikely to be rephrased accurately, degrading correlation with human judgments (see Appendix C.1).
@@ 4.2 Coherent Question Generation
While the above approaches may boost the coherence of PMD rationales, they have limitations. First, since they are training-free, they do not improve the internal coherence of VLMs, rather they only filter and augment their outputs. Further, these interventions take significantly more time and compute due to the need to evaluate the coherence of candidate questions (which requires prompting an NLI model) and generate questions twice (once based on the self-dialog history and once with in-context learning). In practical applications like task guidance, it may be advantageous to use specialized VLMs to improve reliability and speed. [^11] To explore this further, we provide a runtime analysis of various configurations of VLM self-dialog in Appendix E.3. As such, we also explore whether VLMs can be fine-tuned to generate more coherent questions using our automated coherence metrics.
@@ 5 Experimental Results
Evaluated models.
We specifically evaluate InstructBLIP Dai et al. (2023a), LLaVA 1.5-7B Liu et al. (2023), and Llama 3.2-Vision-11B Dubey et al. (2024), small open-source VLMs feasible for online use (important for real-world applications like task guidance). They apply different architectures and training strategies to integrate vision into their LMs. LLaVA and InstructBLIP were not trained on Ego4D, while Llama 3’s training data was not disclosed. For a reference point with state-of-the-art proprietary VLMs, we additionally include results with off-the-shelf GPT-4o OpenAI et al. (2024a). Additional details and prompt templates can be found in Appendix D and E.5.141414In Appendix E.6, we include an additional evaluation of VLMs without generating rationales (as done in prior work).
Human accuracy.
To create a reference point for VLMs’ PMD accuracy and a proxy for data quality and objectiveness, we recruited human annotators for PMD classification. 100 random testing examples were labeled by 3 annotators each, yielding a majority-vote human accuracy of 72.0%. This suggests that PMD itself is fairly subjective and difficult for humans, further necessitating coherent rationalization. More details about this annotation are provided in Appendix C.2.
@@
Table 1: Ego4D-PMD test set results for GPT-4o OpenAI et al. (2024a) and likelihood-based (L) and coherence-based (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
@@ 5.1 Coherent Question Selection Results
Experimental results for coherent question selection interventions are presented in Table 2.151515Hyperparameters and validation results in Appendix E.7. We find that coherence-based re-ranking and in-context learning [^16] We visualize the distribution of question sources (i.e., in-context learning vs. full self-dialog history) in Appendix E.8. both sharply improve the relevance and informativeness in all models, reaching a respective 75.5% and 0.464 bits in LLaVA. This suggests that questions that VLMs find most likely are not naturally the most coherent. Interestingly, accuracy also jumps sharply for InstructBLIP and LLaVA to a maximum of 67.8%, demonstrating that coherent rationales are valuable to accurate PMD, although accuracy still lags slightly behind humans (72.0%). Information gain under these interventions is consistently higher, reaching a maximum of 0.663 bits in LLaVA. This suggests that VLMs can make more confident conclusions given more coherent rationales. [^17] We contextualize these results with a naïve semantic diversity-based ranking baseline in Appendix E.9. Meanwhile, our best configurations outperform GPT-4o in accuracy, relevance, and informativeness at less than 10% of its size Abacha et al. (2025). While GPT-4o takes fewer iterations and exhibits higher information gain (thus making faster and more confident decisions), its rigidity as a closed proprietary VLM prevents further improvements.
@@ 5.2 Coherent Question Generation Results
However, we find that accuracy and informativeness drop slightly from the best results in Table 2 to a respective maximum of 64.2% and 0.340 bits. This suggests that asking more relevant questions is not enough to improve performance globally. While more coherent rationales previously improved accuracy, this instead demonstrates a nontrivial relationship between coherence metrics and task accuracy. We suspect that asking highly relevant questions (which strongly indicate success if answered one way, otherwise a mistake) introduces a trade-off with question difficulty. For elaborate procedures, highly relevant questions may cover multiple aspects or states of a scene. As such, answering these more complex questions may be difficult for VLMs. For example, while the VLM-generated question “Is the soil placed around the seedling with the trowel in the person’s hand?” covers the success conditions of the procedure “Put some soil around the tomato seedling with the gardening trowel in your hand,” answering it requires understanding spatial relations between several objects (i.e., soil, seedling, trowel, and hand). Future work may explore methods to encourage generating simpler questions. [^18] As an initial inquiry here, we experiment with applying an exponential length penalty to VQG in Appendix E.10.
@@ 5.3 Visualizing PMD Performance
An advantage of our automated coherence metrics is the ability to audit VLMs’ reasoning behaviors. In Figure 4, we visualize the distribution of decision error, example-level relevance, and example-level informativeness of four representative approaches. Specifically, we compare vanilla LLaVA (i.e., with likelihood-based question ranking) to variants successively equipped with coherence-based question ranking, in-context learning, and coherence-based DPO fine-tuning. For each example, decision error is calculated by how far the VLM’s success likelihood is from being 100% confident in the correct success label.
Point colors indicate combinations of decision error, relevance, and informativeness, highlighting common local outcomes. Several examples are shown in Figure 5. While cyan points indicate correct decisions with coherent rationales, red points indicate incorrect decisions with incoherent rationales, e.g., asking irrelevant questions about whether a person was wearing protective gear in determining whether a screw had been tightened. Black and indigo points indicate correct decisions with incoherent rationales, while white points indicate incorrect decisions with coherent rationales, suggesting inconsistencies in rationales or failures to interpret them (e.g., correctly identifying a bottle of mustard on the countertop, but then mistakenly detecting it on the floor). Interestingly, green and yellow points with low informativeness typically indicate failures in VQA, e.g., unsure answers, missing the appearance of a trowel, and hallucinating the appearance of a bottle. These examples elucidate the nontrivial relationship between rationale coherence and PMD accuracy observed in earlier results; a variety of errors in both generating and interpreting rationales can cause downstream errors in PMD. We provide an extended discussion of these cases in Appendix E.11.
@@
Example A: Pick up a sink brush from the kitchen slab. Refer to caption Label:    Predicted: Rationale: Is the sink brush… 1. in the person’s hand? Yes
Example E: Fold the cut piece of cloth. Refer to caption Label:    Predicted: Rationale: Is the person… 1. working on a piece of cloth? No
Example B: Tighten the screw. Refer to caption Label:    Predicted: Rationale: Is the person… 1. wearing gloves? No 2. wearing protective gear? No 3. wearing a mask? No
Example F: Put the bolt remover in the lawn tractor. Refer to caption Label:    Predicted: Rationale: Is the bolt remover… 1. in the lawn tractor? Unsure 2. inside the lawn tractor? Unsure 3. attached to the lawn tractor? Unsure
Example C: Paint the stone. Refer to caption Label:    Predicted: Rationale: Is the person… 1. wearing a shirt? Unsure
Example G: Put the trowel in a bin. Refer to caption Label:    Predicted: Rationale: Is the trowel… 1. in a bin? No
Example D: Drop the bottle of mustard on the countertop. Refer to caption Label:    Predicted: Rationale: Is the bottle of mustard… 1. on the countertop? Yes 2. on the floor? Yes
Example H: Put the bottle in the cabinet. Refer to caption Label:    Predicted: Rationale: Is the bottle… 1. in the cabinet? Yes
Figure 5: Sample coherent PMD outputs from LLaVA with coherence-based ranking, representing the range of behaviors observed (as visualized in Figure 4). Images cropped for visibility and space.